calculate the molarity of a solution with a final volume of 500ml prepared by dissolving 25.0ml of methonal (CH3OH density =0.791ml) in choloroform

1
Expert's answer
2022-03-15T11:34:06-0400

Solution:

Liters of CH3OH solution = (500 mL soln) × (1L / 1000 mL) = 0.5 L soln


Density = Mass / Volume

Therefore,

Mass of CH3OH = Density of CH3OH × Volume of CH3OH

Mass of CH3OH = (0.791 g/mL) × (25.0 mL) = 19.775 g


The molar mass of methanol (CH3OH) is 32.04 g/mol

Therefore,

Moles of CH3OH = (19.775 g CH3OH) × (1 mol CH3OH / 32.04 g CH3OH) = 0.6172 mol CH3OH


Molarity = Moles of solute / Liters of solution

Therefore,

Molarity of CH3OH solution = Moles of CH3OH / Liters of CH3OH solution

Molarity of CH3OH solution = (0.6172 mol) / (0.5 L) = 1.2344 mol/L = 1.23 M

Molarity of CH3OH solution = 1.23 M


Answer: The molarity of CH3OH solution is 1.23 M